OpenStudy (anonymous):

Two riders are in a bike race. The second rider starts half an hour late but rides 5 miles per hour faster than the first rider. If they both finish 2 hours after the race starts, how fast are they each riding?

OpenStudy (anonymous):

A. 5 mph and 10 mph B. 12.5 mph and 17.5 mph C. 15 mph and 20 mph D. 10 mph and 15 mph

OpenStudy (mathstudent55):

The both rode the same distance, d miles. Can you try to fill out the following table with the given information. Distance Time Speed First Rider Second Rider

OpenStudy (anonymous):

All I know is time. The first did it in 2 hours and the second did it in an hour and a half.

OpenStudy (mathstudent55):

The distance is the same for both. The two times are known. The speed of the first rider is s. What is the speed of the second rider? Distance Time Speed First Rider d Second Rider d

OpenStudy (anonymous):

s + 5?

OpenStudy (mathstudent55):

Great. Distance Time Speed First Rider d 2 s Second Rider d 1.5 s + 5

OpenStudy (mathstudent55):

The above is what we know. Remember, the distance for both riders is equal. There is one more important piece of information. speed = distance/time Solving for distance, we get distance = speed * time Ok so far?

OpenStudy (anonymous):

Yes, I believe so.

OpenStudy (mathstudent55):

Great. Now using the info of the first rider in the table, what is the distance equal to?

OpenStudy (mathstudent55):

Remember, distance = speed * time Distance is d, replace speed and time with the table values for the first rider.

OpenStudy (anonymous):

2s?

OpenStudy (mathstudent55):

The whole equation is d = 2s. Good. Now do the same using the information of the second rider. What is distance equal to?

OpenStudy (anonymous):

1.5s+5?

OpenStudy (mathstudent55):

I think you meant this: d = 1.5(s+5) which is correct.

OpenStudy (anonymous):

Yes, sorry. Thank you.

OpenStudy (mathstudent55):

1.5 doesn't just multiply s. 1.5 multiplies the quantity s + 5. That's why you need parentheses.

OpenStudy (anonymous):

Ahhh. Alright, I see that now.

OpenStudy (mathstudent55):

Ok. Now we know that d = 2s and d = 1.5(s + 5). d is the same for both equations, so we can write 2s = 1.5(s + 5) Now we have a simple equation with only one variable, s. We can solve for s to find the speed of the first rider.

OpenStudy (mathstudent55):

2s = 1.5(s + 5) First, we distribute the 1.5 on the right side to get rid of the parentheses.

OpenStudy (anonymous):

So would you dived 2s by 1.5?

OpenStudy (mathstudent55):

You can do that, but I think it's easier to distribute the 1.5 on the right side like this: 2s = 1.5s + 1.5 * 5 2s = 1.5s + 7.5

OpenStudy (mathstudent55):

Now subtract 1.5s from both sides.

OpenStudy (anonymous):

So 0.5s=7.5?

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Great. Now, divide both sides by 0.5. What do you get for s?

OpenStudy (anonymous):

s=15? So C?

OpenStudy (mathstudent55):

Correct. s = 15 That means the speed of the first rider is 15 mph. We were told that the second rider's speed is 5 mph higher, so the second rider's speed is 20 mph.

OpenStudy (anonymous):

Awesome! Do you think you can help me with a couple more?

OpenStudy (mathstudent55):

I can do 1 more, but then I need to go. Please start anew post. I'll look for it.